OpenStudy (firejay5):

Find the x and y intercept of the graph of each equation: y = 4x - 2

OpenStudy (lgbasallote):

x -intercept let y = 0 then solve for x y -intercept let x = 0 then solve for y does that help?

OpenStudy (firejay5):

? show me

OpenStudy (lgbasallote):

y = 4x - 2 to get x-intercept let y = 0 0 = 4x - 2 now solve for x

OpenStudy (firejay5):

I am a visual learner, I need to see it

OpenStudy (lgbasallote):

fine...let me do a different example

OpenStudy (firejay5):

thanks

OpenStudy (lgbasallote):

y = 3x + 6 to find the x-intercept..let y = 0 0 = 3x + 6 subtract 6 from both sides -6 = 3x divide both sides by 3 -2 = x so x -intercept = (-2,0) to find the y-intercept let x = 0 y = 3(0) + 6 simplify... y = 6 so the y-intercept is (0,6) got it now?

OpenStudy (firejay5):

to find y intercept for the problem would be (2, -2)

OpenStudy (lgbasallote):

what do you mean (2,-2)?

OpenStudy (firejay5):

what would the y intercept me

OpenStudy (firejay5):

be

OpenStudy (lgbasallote):

let x = 0 y = 4(0) - 2 simplify that..what do you get?

OpenStudy (firejay5):

would it be (0,-2)

OpenStudy (lgbasallote):

yes!!

OpenStudy (firejay5):

okay!!!!